OpenStudy (anonymous):

Don is twice as old as Irene; Ed is 16 years older than Irene. The sum of all three of their ages is 60 years. Find the age of each person. I've calculated the equations with variables and all that. Don's age is D, Irene's is a, and Ed's is E. D= 2a E = 16 + a D + E + a = 60 2i +(16 + a) + a = 60

OpenStudy (anonymous):

I'm going to substitute "i" for "a" because "i" is the symbol for imaginary numbers

OpenStudy (anonymous):

good idea

OpenStudy (anonymous):

D=22 E=27

OpenStudy (anonymous):

D = 2I E = I+16 D+E+I = 60 2I + (I+16) + I = 60 4I + 16 = 60 4I = 44 I = 11 {plug this into the 1st and 2nd} D = 22 E = 27 Hope this helps! :D

OpenStudy (anonymous):

There ya go^ ZhangYan has the right answer! :)

OpenStudy (anonymous):

thanks guys =]